OpenStudy (anonymous):

ok this proportion is driving me nuts... question is the ratio of red to blue marbles is 5 to 7. if there are 156 marble in total, how many red marbles are there?

myininaya (myininaya):

\[\frac{r}{b}=\frac{5}{7}\] \[r+b=156\]

myininaya (myininaya):

do you see how I got those equations?

myininaya (myininaya):

by the way r is red marbles and b is blue marbles

OpenStudy (anonymous):

See out of total 12 (5+7) marbles 5 are red i.e 5/12 so 5 --- * 156 = no of red marbles 12 so no of red marbles is 65

OpenStudy (anonymous):

i do looki at them but was setting up to do cross multiply 5/7=156/x

OpenStudy (anonymous):

5/6 is ratio of red and blue marbles and 156 is total number of marbles so u can't form this relation...

myininaya (myininaya):

\[r=\frac{5b}{7}\] \[\frac{5b}{7}+b=156\] \[\frac{5b+7b}{7}=156\] \[\frac{12b}{7}=156\] \[b=156*\frac{7}{12}=91\] but that means we have r+91=156 which means r=156-91=65